# Large Multimodal Models for Low-Resource Languages: A Survey

### v1

@@ 6 Visual Enhancement Techniques
Visual enhancement techniques aim to improve MT quality by leveraging visual information to provide additional context and resolve ambiguities in the source text. These techniques broadly fall into two main categories: image-guided translation, which uses visual features to enhance the overall translation process, and visual disambiguation, which specifically focuses on resolving ambiguous words or phrases using visual context.
@@
Image-guided translation. A promising direction for improving translation quality for LR languages is the use of image-guided translation approaches. Chowdhury et al. [2018] showed that augmenting neural MT systems with visual features extracted from a pre-trained CNN and integrated into an encoder-decoder architecture can improve translation quality, achieving a bilingual evaluation understudy (BLEU) score of 24.2 for Hindi to English translation. Building upon these ideas, Laskar et al. [2020] developed a multimodal neural MT system with a bidirectional RNN encoder and doubly-attentive decoder for English-Hindi translation. Their system, which combines visual and textual features and employs pre-trained word embeddings from monolingual data, outperforms a text-only baseline, achieving a BLEU score of 33.57 versus 27.75 on the test set.
@@
Subsequent studies Shi and Yu (2022); Meetei et al. (2023a, b); Haq et al. (2024) have demonstrated the effective use of visual information for improving MT in LR settings, particularly for the English-Hindi language pair. Meetei et al. [2023a] proposed a video-guided multimodal MT framework that incorporates spatio-temporal video features, showing improvements of up to +4.2 BLEU over text-only baselines for English to Hindi translation, while Meetei et al. [2023b] explored multimodal translation for news domain data, showing that ResNet-based image features outperform VGG-based features and improve BLEU scores by +1.8 points. Additionally, Shi and Yu [2022] explored different approaches for extracting and integrating image features using VGG and ResNet models, achieving a +3 BLEU improvement over text-only translation. More recently, Haq et al. [2024] presented a context-aware transformer model that integrates visual features via BERT encoding, demonstrating consistent improvements over text-only baselines. Across all studies, qualitative analyses confirmed that visual cues are particularly beneficial for handling rare words and domain-specific terms, with both image and video modalities helping to resolve ambiguity and improve translation quality in LR scenarios.
@@
Visual disambiguation. While image-guided translation aims to enhance overall translation quality by integrating visual context, the visual disambiguation techniques focus on task-specific ambiguities by grounding them in visual information. In this regard, studies revolving around the creation of Visual Genome datasets for LR languages, such as Hindi Parida et al. (2019), Bengali Sen et al. (2022) and Hausa Abdulmumin et al. (2022), have played a pivotal role in advancing visual disambiguation techniques.
@@
Another work highlighting the benefits of using visual features for disambiguation is presented by Jain et al. [2021]. Their model, called MURAL, shows strong performance on text-to-image retrieval tasks, where it manages to retrieve relevant images for ambiguous queries. This finding is also supported by the qualitative examples, where MURAL successfully disambiguates word senses based on visual context. In addition, Kovath et al. [2024] proposed a co-attention mechanism for Malayalam VQA that allows the model to jointly learn attention over both textual and visual inputs, demonstrating improved performance over baseline approaches that use only textual features.
@@
Comparative analysis of visual enhancement techniques. Image-guided translation consistently demonstrates performance improvements over text-only baselines for LR languages, though effectiveness varies with language pair, dataset size, and translation direction. These approaches excel at handling semantic ambiguities and culturally-specific concepts, but their success depends heavily on the quality of extracted visual features. A key limitation is the reliance on high-quality image-text pairs, which are often scarce for LR languages. While these techniques improve translation quality, they also introduce computational overheads. Future work should focus on developing more efficient visual feature extraction methods and better approaches for leveraging visual information with limited paired data.
@@ 7 Cross-Modal Transfer Learning
Modality transfer. Modality transfer addresses the challenge of transferring knowledge between different modalities to improve performance on LR tasks. A diversity of approaches have been used to achieve modality transfer. Chen et al. [2023] proposed a progressive transfer learning strategy that leverages both general pre-training (Kinetics-400 for visual and CC25 for language) and domain-specific pre-training (sign-to-gloss translation) to bridge modalities for sign language translation. Amalas et al. [2024] introduced a data-driven approach for selecting source languages and demonstrated that multilingual pre-training outperforms monolingual pre-training for text-to-speech systems. Yeo et al. [2024] tackled LR visual speech recognition by using Whisper’s automatic transcriptions to generate training labels from unlabeled multilingual audio-visual data. For Arabic handwriting recognition, Bhatia et al. [2024] employed modality transfer through an architecture combining SwinV2 for visual encoding and RoBERTa for text decoding, while Tran and Thanh [2024] demonstrated successful modality transfer for Vietnamese through extensive pre-training of both vision and language components, combined with automated data curation methods. Notably, Onuoha and Uba [2024] challenged the assumptions about multimodal integration through their study of Igbo minimal pairs. Their findings show that native Igbo speakers can accurately distinguish minimal pairs through audio alone, suggesting that the benefits of cross-modal integration may be more relevant for non-native speakers than fluent ones.
@@
Language transfer. Language transfer is an approach for leveraging knowledge from resource-rich languages to improve model performance on LR languages. Recent work demonstrates several effective strategies. For example, Wang et al. [2023b] adapted MDETR to new languages by using adapters and code-switching without relying on MT data. Kim et al. [2023] focused on learning general speech knowledge from English for lip reading, and combining it with language-specific audio features. dos Santos et al. [2023] proposed to use data augmentation and contrastive learning to improve multilingual CLIP models for LR languages. Nortje et al. [2024] showed that initializing a Yorùbá few-shot word learning model with weights from an English speech-image model substantially improves performance. These approaches share the common theme of transferring learned representations and knowledge from high-resource languages (typically English), while developing techniques to efficiently adapt and fine-tune models for target LR languages.
@@ 8 Architectural Innovations
Some recent architectural innovations in the context of LR languages have focused on adapting the CLIP architecture, originally introduced by Radford et al. [2021]. One such example is the LowCLIP Asgarov and Rustamov (2024) model, which replaces the original text encoder trained primarily on English text with a multilingual BERT (mBERT). The authors evaluated various lightweight image encoders, such as EfficientNet-B0 and Tiny Swin Transformer, for a more computationally efficient approach, while also targeting LR languages like Azerbaijani. To compensate for the lighter architecture and the scarcity of image-text pairs in Azerbaijani, LowCLIP leverages synthetic data generation via MT for text features, and image augmentation techniques, such as crop and rotation, for image features. In contrast, XtremeCLIP Tang et al. (2023) took a different approach, where the authors introduced a parameter-efficient method that only tunes a small prototype matrix, while keeping the visual and text encoders frozen. Their model also leverages contrastive learning to provide additional supervised signals in LR settings. Collectively, these efforts extend the applicability of CLIP to multimodal image retrieval tasks.
Another approach for multimodality in the context of LR languages is introduced by Wu et al. [2019]. The approach combines two existing methods, a translation-based one and an alignment-based one, into a unified architecture for improving image captioning. The framework employs a model that first generates high-quality English captions, which are then used together with the images to produce captions in the LR language. The model achieves a fine-grained alignment between visual elements and captions in both languages via a cycle-consistency constraint, outperforming existing methods on standard metrics.
@@
Together, the architectural innovations introduced by LowCLIP, XtremeCLIP, FEWVLM and similar frameworks demonstrate various viable alternatives for making multimodal models more accessible for LR languages.

### v2

@@ 6 Visual Enhancement Techniques
Visual enhancement techniques aim to improve MT quality by leveraging visual information to provide additional context and resolve ambiguities in the source text. These techniques broadly fall into two main categories: image-guided translation, which uses visual features to enhance the overall translation process, and visual disambiguation, which specifically focuses on resolving ambiguous words/phrases via visual context.
@@
Image-guided translation. A promising direction for improving translation quality for LR languages is the use of image-guided translation approaches. Dutta Chowdhury et al. [2018-dutta] showed that augmenting neural MT systems with visual features extracted from a pre-trained CNN and integrated into an encoder-decoder architecture can improve translation quality, achieving a bilingual evaluation understudy (BLEU) score of 24.2 for Hindi to English translation. Building upon these ideas, Laskar et al. [2020-laskar, 2021-laskar-hindi] developed a multimodal neural MT system with a bidirectional RNN encoder and a doubly-attentive decoder for English-Hindi translation. Their system, which combines visual and textual features, and employs pre-trained word embeddings from monolingual data, outperforms a text-only baseline, achieving a BLEU score of 33.57 versus 27.75 on the test set.
@@
Subsequent studies [2023-gain, 2022-shi, 2023-meetei-cues, 2023-meetei-hindi, 2024-haq] have demonstrated the effective use of visual information for improving MT in LR settings, particularly for the English-Hindi language pair. Meetei et al. [2023-meetei-cues] proposed a video-guided multimodal MT framework that incorporates spatio-temporal video features, showing improvements of up to +4.2 BLEU over text-only baselines for English to Hindi translation, while Meetei et al. [2023-meetei-hindi] explored multimodal translation for news domain data, showing that ResNet-based image features outperform VGG-based features and improve BLEU scores by +1.8 points. Additionally, Shi et al. [2022-shi] explored different approaches for extracting and integrating image features using VGG and ResNet models, achieving a +3 BLEU improvement over text-only translation. Another contribution is presented by Gain et al. [2023-gain], who showed how visual context enhances translation robustness under noisy conditions (e.g. OCR errors), even when image relevance is reduced. Extending this line of work, Tayir et al. [2024-tayir] demonstrated that visual context can bridge structural gaps in distant language pairs, such as English-Uyghur, by introducing a visual masked language modeling approach for unsupervised multimodal MT. Similarly, Tayir et al. [2025-tayir] improved translation for the same language pairs by harnessing varying-granularity image features in low-resource settings.
More recently, Haq et al. [2024-haq] presented a context-aware transformer model that integrates visual features via BERT encoding, demonstrating consistent improvements over text-only baselines. In a related direction, Lekshmy et al. [2022-h-o] developed an English-Malayalam vision-aided translation system for visually impaired users, employing multimodal machine learning techniques to perform object recognition and generate translated descriptions in real-time.
Across all studies, qualitative analyses confirmed that visual cues are particularly beneficial for handling rare words and domain-specific terms, with both image and video modalities helping to resolve ambiguity and improve translation quality in LR scenarios.
@@
Visual disambiguation. While image-guided translation aims to enhance overall translation quality by integrating visual context, the visual disambiguation techniques focus on task-specific ambiguities by grounding them in visual information. In this regard, studies revolving around the creation of Visual Genome datasets for LR languages, such as Hindi [2019-parida], Bengali [2022-sen] and Hausa [2022-abdulmumin], have played a pivotal role in advancing visual disambiguation techniques. Building upon previous work, Parida et al. [2021-parida] explored this line of research by developing a multimodal NMT system for English-Bengali using object tags extracted from images as auxiliary input, while Nortje et al. [2024-nortje-2] introduced an innovative few-shot learning approach for visually-prompted keyword localization in Yoruba.
@@
Several studies have investigated the use of visual features for disambiguation [2021-jain, 2022-jian, 2024-kovath, 2020-laskar, 2022-laskar, 2021-laskar-english-assamese]. For example, Jain et al. [2021-jain] highlighted the benefits of using visual features for disambiguation. Their model, called MURAL, shows strong performance on text-to-image retrieval tasks, where it manages to retrieve relevant images for ambiguous queries. This finding is also supported by the qualitative examples, where MURAL successfully disambiguates word senses based on visual context. In addition, Kovath et al. [2024-kovath] proposed a co-attention mechanism for Malayalam VQA that allows the model to jointly learn attention over both textual and visual inputs, demonstrating improved performance over baselines using only textual features.
@@
Comparative analysis of visual enhancement techniques. Image-guided translation consistently demonstrates performance improvements over text-only baselines for LR languages, though effectiveness varies with dataset size and translation direction. These approaches excel at handling semantic ambiguities and culturally-specific concepts, but their success depends heavily on the quality of extracted visual features. A key limitation is the reliance on high-quality image-text pairs, which are often scarce for LR languages. While these techniques improve translation quality, they also introduce computational overheads. Future work should focus on developing more efficient visual feature extraction methods and better approaches for leveraging visual information with limited paired data.
@@ 7 Cross-Modal Transfer Learning
Cross-modal transfer learning represents a critical approach for LR languages, allowing models to harness knowledge from data-rich modalities or languages to improve performance in resource-constrained settings. Unlike traditional transfer learning, which operates within a single modality, cross-modal transfer must bridge the significant gap between different types of data representations. This is conceptually similar to how a person might use their understanding of written language to help learn a sign language, or how knowledge of one spoken language can facilitate learning another. In the context of low-resource languages, two primary transfer directions have emerged: modality transfer, which moves knowledge between different data types (e.g. from text to speech), and language transfer, which leverages high-resource languages to improve performance in low-resource ones.
@@
Modality transfer. Modality transfer addresses the challenge of transferring knowledge between different modalities to improve performance on low-resource tasks. This approach is particularly valuable when certain modalities have more abundant data than others. For example, text data is often easier to collect than paired speech data for many languages. The fundamental challenge lies in bridging the representational gap between modalities, since text operates in a discrete symbolic space, while speech and vision exist in continuous signal spaces with very different statistical properties. Successful modality transfer requires finding meaningful mappings between these different representational spaces. A diversity of approaches has been used to achieve modality transfer. Chen et al. [2023-chen] proposed a progressive transfer learning strategy that leverages both general pre-training (Kinetics-400 for visual and CC25 for language) and domain-specific pre-training (sign-to-gloss translation) to bridge modalities for sign language translation. Amalas et al. [2024-amalas] introduced a data-driven approach to select source languages and demonstrated that multilingual pre-training outperforms monolingual pre-training for text-to-speech systems. Wu et al. [wu2022pairs] developed a captioning approach via multi-objective optimization that addresses the challenge of utilizing both triplet datasets (image, HR language, LR language) and large-scale paired datasets during training. Yeo et al. [2024-yeo] tackled LR visual speech recognition by using Whisper-based automatic transcriptions to generate training labels from unlabeled multilingual audio-visual data. For Arabic handwriting recognition, Bhatia et al. [bhatia2024qalammultimodalllm] employed modality transfer through an architecture combining SwinV2 for visual encoding and RoBERTa for text decoding, while Tran et al. [2024-tran-lavy] demonstrated successful modality transfer for Vietnamese through extensive pre-training of both vision and language components, combined with automated data curation methods. Notably, Onuoha et al. [2024-onuoha] challenged the assumptions about multimodal integration through their study of Igbo minimal pairs. Their findings show that native Igbo speakers can accurately distinguish minimal pairs through audio alone, suggesting that the benefits of cross-modal integration may be more relevant for non-native speakers.
Table 6: Overview of architectural innovations for low-resource multimodal learning. V = Vision, T = Text, A = Audio. Although Cycle-Attn is evaluated on EN and DE (high-resource), it is included as a key methodological reference. The authors simulated a low-resource scenario using the limited Multi30K dataset to demonstrate the efficacy of knowledge transfer from a rich monolingual corpus (EN) via cycle consistency constraints.
Model/Method Year Languages Modalities Task Approach Category Key Innovation Cycle-Attn [2019-wu] 2019 EN, DE V, T Image Captioning Translation+Alignment Cycle consistency constraint for cross-lingual alignment Multi-task Adversarial [2022-mamta] 2022 EN, HI T, A Sentiment Analysis Adversarial Learning Cross-lingual transfer via shared embeddings FEWVLM [jin2022good] 2022 EN, HI V, T VL Understanding Prompt Engineering Few-shot prompting with moderate-size VLM Amharic Captioning [solomon2023amharic] 2023 Amharic V, T Image Captioning Attention-based DNN Visual attention + Bi-GRU decoder Auxiliary CTC [2023-chen] 2023 102 langs A, T Multilingual ASR CTC Conditioning LID-conditioned auxiliary objectives Sanskrit-Malayalam NMT [rahul2023morphology] 2022 SA, ML T, A Machine Translation Multimodal NMT Morphology + WSD embedding fusion XtremeCLIP [2023-tang-xtremeclip] 2023 EN, HI V, T VL Understanding Parameter-efficient Prototype affinity matching (5-7K params) LowCLIP [2023-Asgarov] 2024 Azerbaijani V, T Image Retrieval Efficiency-first mBERT + lightweight image encoders Yoruba ASR [rahmon2024speech] 2024 EN, YO T, A Speech Recognition Transfer Learning MFCC-based acoustic modeling Llama 3 [lamma] 2024 200 langs V, T General Multimodal Foundation Model Native multimodal MoE architecture DeepSeek-V3 [DeepSeek] 2024 14 langs V, T General Multimodal MoE Architecture MLA + FP8 training efficiency Claude 4 [anthropic2025claude] 2025 15 langs V, T General Multimodal Foundation Model Dual-mode reasoning operation Apple AFM [DBLP:journals/corr/abs-2407-21075] 2024 16 langs V, T On-device/Server Distillation+QAT 2-bit quantization for edge deployment MMaDA [MMaDA] 2025 60+ langs V, T Multimodal Diffusion Unified Diffusion Discrete diffusion language modeling MixLoRA [Mixture_of_LoRA] 2024 15 langs V, T Instruction Tuning Dynamic PEFT Conditional mixture routing for adaptation
@@
Language transfer. Language transfer is an approach to harness knowledge from HR languages to improve model performance on LR languages. Recent work demonstrated several effective strategies. For instance, Wang et al. [2023-wang-adapting] adapted MDETR to new languages by using adapters and code-switching without relying on MT data. Cheema et al. [cheema2024adapting] presented ViLanOCR, a novel approach that adapts multilingual vision language transformers for low-resource Urdu optical character recognition by leveraging the Swin encoder and mBART-50 decoder. Kim et al. [2023-kim] focused on learning general speech knowledge from English for lip reading, and combining it with language-specific audio features. Aruna Gladys et al. [vetriselvi2024sentiment] proposed a multimodal representation learning framework that uses cross-lingual transfer learning to analyze sentiment in LR language datasets, demonstrating significant performance improvements for Tamil language sentiment analysis. Chen et al. [2023-wchen] improved multilingual ASR by conditioning models on language identity predictions from early layers to enhance performance across numerous languages. dos Santos et al. [2023-santos] proposed to use data augmentation and contrastive learning to improve multilingual contrastive language-image pre-training (CLIP) models for LR languages. Nortje et al. [2024-nortje] showed that initializing a Yoruba few-shot word learning model with weights from an English speech-image model substantially improves performance. These approaches share the common theme of transferring learned representations and knowledge from HR languages (typically English), while developing techniques to adapt and fine-tune models for target LR languages.
Table 7: Computational requirements for architectural innovations. A dash line indicates that the respective information was not reported in the original paper.
[…]
The effectiveness of language transfer methods varies significantly based on linguistic similarity, writing systems, and cultural context. For instance, transfer between closely related languages (such as Spanish to Portuguese) typically outperforms transfer between distant language families (such as English to Tamil). The methods described above demonstrated different approaches to this challenge: Wang et al. [2023-wang-adapting] focused on architecture adaptation through adapters, while Kim et al. [2023-kim] emphasized feature-level knowledge transfer. Meanwhile, Nortje et al. [2024-nortje] showed that even initialization from a different language can provide substantial benefits. For practitioners working with specific low-resource languages, the choice between these approaches should consider both linguistic factors and computational constraints.
@@ 8 Architectural Innovations
Table 8: Performance metrics for low-resource multimodal architectures. All values extracted directly from source papers. Baseline methods and improvement calculations are specified for reproducibility. Full FT = full fine-tuning; Aug. = augmentation; – = not applicable or not reported.
@@
Architectural innovations for low-resource multimodal learning focus on designing model structures that can effectively leverage limited data while maintaining reasonable computational requirements. The fundamental challenge lies in balancing model capacity (ability to learn complex patterns) with sample efficiency (ability to learn from limited examples). While simply scaling down large models designed for high-resource settings is one approach, the most successful innovations in this space incorporate architectural elements specifically designed to address the constraints of low-resource scenarios. These innovations generally fall into three categories: (1) efficiency-focused adaptations of existing architectures, (2) parameter-efficient fine-tuning methods, and (3) novel architectures designed specifically for low-resource multimodal learning. In Table 6, we provide a systematic overview of these architectural innovations, categorized by approach type, supported modalities, and target tasks. Tables 7 and 8 complement this overview with quantitative analyses of computational requirements and empirical performance, enabling direct comparison across methods with varying resource constraints.
Some recent architectural innovations in the context of LR languages have focused on adapting the CLIP architecture [radford2021learning]. One such example is the LowCLIP model [2023-Asgarov], which replaces the original text encoder trained primarily on English text with a multilingual BERT (mBERT). The authors evaluated various lightweight image encoders, such as EfficientNet-B0 and Tiny Swin Transformer, for a more computationally efficient approach, while also targeting LR languages like Azerbaijani. To compensate for the lighter architecture and the scarcity of image-text pairs in Azerbaijani, LowCLIP leveraged synthetic data generation via MT for text features, and image augmentation techniques, such as crop and rotation, for image features. In contrast, XtremeCLIP [2023-tang-xtremeclip] took a different approach, in which the authors introduced a parameter-efficient method that only tunes a small prototype matrix, while keeping the visual and text encoders frozen. Their model also employs contrastive learning to provide additional supervisory signals in LR settings. Collectively, these efforts extend the applicability of CLIP to multimodal image retrieval tasks.
Table 9: Image encoder performance comparison for low-resource image retrieval. Results are taken from LowCLIP [2023-Asgarov].
Image Encoder Params GFLOPs Size (MB) mAP COCO Flickr8k Flickr30k ResNet-50 25.6M 4.09 97.8 0.80 0.76 0.73 EfficientNet-B0 5.29M 0.39 20.5 0.81 0.85 0.87 ViT-Base 86.6M 17.56 330.3 0.71 0.80 0.70 Swin-Tiny 28.3M 4.49 108.2 0.80 0.84 0.79
[…]
Another approach for multimodality in the context of LR languages is introduced by Wu et al. [2019-wu]. The approach combines two existing methods, a translation-based one and an alignment-based one, into a unified architecture to improve image captioning. The framework employs a model that first generates high-quality English captions, which are then used together with the images to produce captions in the LR language. The model achieves a fine-grained alignment between visual elements and captions in both languages via a cycle-consistency constraint, outperforming existing methods on standard metrics.
Table 10: Comparison of parameter-efficient fine-tuning methods for low-resource vision-language understanding. VE = Visual Entailment, VQA = Visual Question Answering, IC = Image Classification. Results are taken from XtremeCLIP [2023-tang-xtremeclip].
Method Trainable Params VE VQA IC Avg. Training Time Zero-shot 0 33.74 52.03 39.17 42.89 Full fine-tuning 149M 51.10 54.10 28.14 51.12 hours LLRD 149M 57.23 53.88 31.36 53.60 hours BitFit 176-178K 59.56 54.72 41.61 55.66 minutes BiNor 208-210K 59.54 54.75 41.73 55.67 minutes CLIP-Adapter 131-262K 59.21 54.21 44.88 55.45 minutes Tip-Adapter 5-10M 59.67 54.70 45.12 55.62 minutes XtremeCLIP 5-7K 62.06 59.21 48.30 57.73 20 minutes LoRA ($r=4$) $\approx$4K 65.39 minutes LoRA ($r=16$) $\approx$16K 65.50 minutes MixLoRA ($E=16$) $\approx$8K/layer 67.17 hours
[…]
Foundation models represent a qualitatively different design point, prioritizing broad capability over resource efficiency. We include them here to contextualize the computational differences that shape research accessibility. Dubey et al. [lamma] introduced the Llama 3 series with models ranging from 8B to 405B parameters, officially supporting 8 languages (English, German, French, Italian, Portuguese, Hindi, Spanish, and Thai), with experimental multilingual capabilities on a broader set via the speech interface (34 languages). Llama 3 multimodal extensions for image, video, and speech understanding were described in their technical report, but remain under development and have not been publicly released along with the paper. In a similar endeavor, Liu et al. [DeepSeek] presented DeepSeek-V3, a 671B-parameter MoE language model (37B active parameters per token) with multi-head latent attention (MLA) and FP8 mixed-precision training. It is important to note that DeepSeek-V3 is a text-only language model without native vision or audio capabilities. However, we include it to put efficient training strategies into perspective (DeepSeek-V3 requires 2.788M H800 GPU-hours, costing approximately \$5.6M) and better inform future multimodal model development. For edge deployment scenarios, Gunter et al. [DBLP:journals/corr/abs-2407-21075] introduced Apple Intelligence Foundation Models with a novel Parallel-Track MoE architecture optimized for on-device processing, supporting 16 languages with 2-bit quantization-aware training. The prevalence of MoE architectures in these recent developments demonstrates the effectiveness of expert-based scaling for multimodal tasks, as also observed by Mu et al. [DBLP:journals/corr/abs-2503-07137]. Additionally, Yang et al. [MMaDA] proposed a unified diffusion architecture that combines multimodal understanding with generation capabilities, offering new perspectives on architectural design for LR contexts.
Table 11: Design strategies and trade-offs in multimodal architectures for low-resource settings. Core strategies are grouped by methodological approach.
[…]
For specific language families and modality combinations, several innovative architectures have been proposed. Solomon et al. [solomon2023amharic] developed a hybridized attention-based deep neural network for Amharic language image captioning, combining a CNN encoder with visual attention mechanisms and a bidirectional GRU decoder, achieving significant improvements in terms of BLEU. Rahul et al. [rahul2023morphology] introduced a multimodal neural machine translation system between Sanskrit and Malayalam, which embeds morphology and word sense disambiguation awareness. It utilizes both textual and speech modalities via a two-level fusion approach of transform-based feature vectors. For African languages, Rahmon et al. [rahmon2024speech] presented a speech recognition model for Yoruba that employs acoustic and language modeling with sequential MFCC features, achieving 83% accuracy in speech-to-text conversion. For sentiment analysis, Mamta et al. [2022-mamta] explored multilingual, multi-task and adversarial learning approaches to transfer knowledge from HR languages to LR scenarios, leveraging shared semantic spaces through cross-lingual word embeddings. For Arabic, Alwajih et al. [2024-Alwajih] introduced Peacock, a comprehensive family of LMMs with strong vision and language capabilities, alongside Henna, a benchmark for evaluating culturally-aware Arabic LMMs, further helping to bridge the gap between high-resource and low-resource languages, while addressing unique linguistic and cultural characteristics.
In Table 11, we synthesize the design trade-offs across architectural approaches, organized by methodological strategy. Three principal patterns emerge from our analysis. First, parameter-efficient adaptation methods (XtremeCLIP, LowCLIP, FEWVLM, MixLoRA) achieve competitive performance, while reducing trainable parameters by 3-5 orders of magnitude compared with full fine-tuning, making the corresponding models accessible to researchers with limited computational resources. Second, cross-lingual transfer approaches (Cycle-Attn, Amharic Captioning) effectively leverage high-resource language supervision, typically English, to bootstrap performance in target languages. However, this creates structural dependency on pivot language quality and availability. Third, foundation models occupy a distinct design regime. Since Llama 3 requires $3.8\times 10^{25}$ FLOPs and DeepSeek-V3 consumes 2.788M H800 GPU-hours for pre-training, these models remain inaccessible to most research groups focused on low-resource languages. The practical implication is that parameter-efficient methods currently offer the most viable path for researchers operating under resource constraints, while foundation models may serve as upstream components for transfer learning when API access or pre-trained weights are available.
The computational requirements documented in Table 7 reveal a structural divide with sociolinguistic implications. While parameter-efficient methods like XtremeCLIP (5-7K parameters, 20 minutes on one GPU) remain accessible, foundation models require resource-intensive infrastructure (Llama 3 consumes $3.8\times 10^{25}$ FLOPs across 16K H100 GPUs; DeepSeek-V3 requires 2.79M H800 GPU hours with an estimated cost of \$5.6M). This asymmetry matters because LLMs exhibit systematic bias in knowledge acquisition. Indeed, new knowledge is learned less efficiently in LR languages, transfers less effectively to them, and is overwritten more easily by HR language information [wu2025inequalities]. The implication is that scaling alone is not sufficient to achieve equity. Therefore, architectural innovations must explicitly counteract these biases.
Federated learning offers a technical framework aligned with data sovereignty principles, enabling collaborative training without data centralization [lin2022fednlp]. Recent work demonstrates feasibility for multilingual LR settings. For example, federated prompt tuning achieves competitive performance while preserving data locality [chen2025breaking], and differential privacy integration protects against gradient inversion attacks [yang2025dpfpl]. For multimodal LR applications, federated approaches could enable geographically-distributed language communities to collaboratively improve models without ceding control over culturally-sensitive audiovisual data.
9 Evaluation Challenges
Evaluation remains one of the most underdeveloped aspects of research on LMMs for LR languages. While the field has made significant strides in dataset creation, fusion strategies, and architectural innovations, the ways for measuring success have not kept pace. The lack of consistent and culturally-grounded evaluation protocols severely hampers the ability of researchers to compare models, reproduce findings, or interpret results in real-world contexts.
Limitations of standard metrics across cultural contexts. Most evaluation pipelines for LR multimodal models rely on automatic metrics originally designed for high-resource and predominantly Western-centric settings. Metrics such as BLEU, ROUGE, accuracy, and F1 implicitly assume that reference annotations reflect shared cultural, visual, and linguistic grounding. This assumption frequently fails in low-resource contexts.
One such case can be observed in multimodal tasks such as visual question answering, image captioning, and meme understanding, where the visual content itself often encodes culturally-specific assumptions regarding object salience, social roles, or everyday activities. For instance, a model trained primarily on Western image datasets may fail to recognize culturally significant objects (e.g. traditional clothing, local foods, religious symbols, etc.) that are common in LR language contexts. When benchmarks are translated or minimally adapted from high-resource languages, models may achieve high lexical overlap with reference answers while still producing outputs that are culturally inappropriate, semantically misleading, or pragmatically implausible for native speakers. These issues are further exacerbated in knowledge-intensive evaluations derived from English-centric benchmarks. For example, questions about local festivals, historical events, or social customs require cultural context that translation alone cannot provide. As a result, standard metrics may overestimate progress or mask systematic failures that are only visible through culturally-grounded evaluation.
Dataset heterogeneity and comparability issues. A second major challenge concerns dataset heterogeneity, as existing studies evaluate multimodal models on datasets with widely distinct characteristics and assumptions. Many studies rely on translated versions of high-resource benchmarks, such as extensions of Multi30K for Ukrainian [2023-saichyshyna] or Visual Genome variants for Bengali [2022-sen], Hausa [2022-abdulmumin], and Hindi [2019-parida]. While translation-based approaches enable rapid benchmark construction, they often introduce Western cultural biases and may fail to reflect authentic language use or visual grounding in target communities. In contrast, newly introduced language-specific datasets, such as DravidianMultiModality [2021-Chakravarthi], RoMemes [Pais2024RoMemesAM], and ArabSign [2022-hamzah], better capture genuine linguistic and cultural phenomena, but typically suffer from limited coverage, non-standardized annotation protocols, and heterogeneous quality control practices, making cross-study comparison difficult. As a result, performance improvements reported across such heterogeneous evaluation settings are often not directly comparable.
Recommendations for fair evaluation practices. Based on our analysis, we propose the following recommendations for evaluation in LR multimodal research:
  • 1.
    Report multiple metrics. Studies should report multiple complementary metrics that capture different aspects of performance. In the context of machine translation tasks, researchers should report BLEU alongside COMET, or human evaluation scores. Another example in the context of VQA tasks, exact-match accuracy should be accompanied by relaxed matching that accounts for morphological variants and, when possible, human judgment of answer correctness.
  • 2.
    Perform culturally-grounded human evaluation. In addition to a diverse set of automated metrics, we believe that human evaluation conducted by native speakers from the target language community also plays a crucial role. Evaluators should assess whether outputs sound natural to native speakers, whether they are culturally appropriate, whether they convey the intended meaning accurately, and (for VQA) whether answers are semantically correct, even if worded differently from the reference.
  • 3.
    Develop and use standardized benchmarks. The field needs publicly available test sets for LR multimodal evaluation, following examples like SEACrowd [2024-lovenia] for Southeast Asian languages and CreoleVal [lent-etal-2024-creoleval] for Creole languages. Such benchmarks should cover different task types (VQA, captioning, translation, classification), include culturally-accurate content created together with language communities, provide multiple correct answers to account for natural variation, and document how data was labeled.
  • 4.
    Compare to sensible baselines. Rather than reporting absolute performance in isolation, studies should contextualize results relative to unimodal baselines (text-only or vision-only) to demonstrate the benefits of multimodal approaches, random and majority-class baselines to establish task difficulty, prior work on the same dataset when available, and performance on related HR languages to quantify the LR gap.
As shown above, evaluation challenges remain a major problem in LR multimodal research, but some steps have already been taken towards fixing this gap. Although standard metrics represent a great starting point for evaluation, they are designed for English and often miss what matters for LR languages and their cultural contexts. Solving these challenges requires creating culturally-appropriate benchmarks, using multiple and diverse evaluation metrics, as well as involving language communities in the evaluation process.